matt needs 45 liters of 20% solution. He has only 15% solution and 30% solution on hand. How many liters of each solution should he combine to make the mixture?
Answer by CharlesG2(834)

matt needs 45 liters of 20% solution. He has only 15% solution and 30% solution on hand. How many liters of each solution should he combine to make the mixture?
% in A * amount A + % in B * amount B = % in C * amount C
call amount A x, and amount B 45-x
15% * x + 30% * (45 - x) = 20% * 45
0.15x + 0.3 * 45 - 0.3x = 0.2*45
0.15x + 13.5 - 0.3x = 9
-0.15x = -4.5 (divide both sides by -1)
0.15x = 4.5
x = 30 liters of A (the 15% solution)
and 15 liters of B (the 30% solution)